OpenStudy (anonymous):

1-x xpower3-1 xpower3 +8 ---- divided by -------- x--------- x+2 x+5 x+5

OpenStudy (anonymous):

\[\frac{\frac{1-x}{x+2}}{\frac{x^3-1}{x+5} * \frac{x^3+8}{x+5}}\] That it?

OpenStudy (anonymous):

Their answer is -(xpower2-2x+4) ------------- xpower2+x+1 HOW DID THEY GET THIS ANSWER??????

OpenStudy (anonymous):

Was that the right question as I wrote it?